# Tight frames over the quaternions and equiangular lines

### v2

@@ 5 Projective unitary equivalence
This leads to a “linear system” $C(W^{*}W)=(V^{*}V)C$ in the scalars $\alpha_{j}$. However, due to the noncommutativity of the quaternions, this can not be solved by Gauss elimination, unless one first converts it to a linear system over $\mathbb{R}$ (in the coordinates of the $\alpha_{j}$). What is usually done in the real and complex cases is to consider a collection of invariants: the $m$-products, which completely characterise projective unitary equivalence [CW16]. We now look at the analogue of these (also see [KMW19] for subspaces of $\mathbb{C}^{d}$).
@@
$\Delta\bigl{(}(U$ $v_{j_{1}})\alpha_{j_{1}},(Uv_{j_{2}})\alpha_{j2},\ldots,(Uv_{j_{m}})\alpha_{j_{m}}\bigr{)}$
@@
$=\overline{\alpha_{j_{1}}}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})\alpha_{j_{1}},$
the $m$-products are projective unitary invariants of $(v_{j})$ up to congruence, and real frames are characterised by having real $m$-products. A quaternion $q$ is determined up to congruence by its real part $\mathop{\rm Re}\nolimits(q)$ and its norm $|q|$, and so we can define (reduced) $m$-products as a pair of real numbers
@@
These are projective unitary invariants. For the complex case, the $m$-products are projective unitary invariants, which depend only the cycle $(j_{1},\ldots,j_{m})$, and a small set of $m$-products corresponding to a basis for the cycle space of the frame graph of $(v_{j})$ provide a set of invariants which characterise projective unitary equivalence (see [CW16]). We can not yet make a similar claim in the quaternionic case, though we do imagine that the $m$-products do characterise projective unitary equivalence.
The dependence of $m$-products on only the associated $m$-cycle in the frame graph does follow, by the calculation
$a\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})a^{-1}=\Delta(v_{j_{2}},v_{j_{3}},\ldots,v_{j_{m}},v_{j_{1}}),\qquad a={\langle v_{j_{1}},v_{j_{2}}\rangle\over|\langle v_{j_{1}},v_{j_{2}}\rangle|},$
@@
$\qquad=|\langle v_{1},v_{2}\rangle|^{2}|\langle v_{2},v_{3}\rangle|^{2}\cdots|\langle v_{k-1}v_{k}\rangle|^{2}\Delta(v_{k},v_{k+1},\ldots,v_{m},v_{1},w_{1},w_{2},\ldots,w_{n}).$
@@
$\langle v_{k+1},v_{k}\rangle\langle v_{k+2},v_{k+1}\rangle\cdots\langle v_{m},v_{m-1}\rangle\langle v_{1},v_{m}\rangle\langle v_{2},v_{1}\rangle\cdots\langle v_{k-1},v_{k-2}\rangle\langle v_{k},v_{k-1}\rangle$
×$\qquad\times\langle v_{k-1},v_{k}\rangle\langle v_{k-2},v_{k-1}\rangle\cdots\langle v_{1},v_{2}\rangle\langle w_{1},v_{1}\rangle\langle w_{2},w_{1}\rangle\cdots\langle w_{n},w_{n-1}\rangle\langle v_{k},w_{n}\rangle,$
which simplifies to the right hand side, since $\langle v_{j+1},v_{j}\rangle\langle v_{j},v_{j+1}\rangle=|\langle v_{j},v_{j+1}\rangle|^{2}\in\mathbb{R}$ commutes with any quaternion.
@@
and the permutations of the vectors which preserve these $5$-products is indeed $A_{6}$. With the present theory, this does not yet establish that $A_{6}$ is the projective symmetry group.
@@
Once the unit scalars $\alpha_{j}$ corresponding to a basis $[v_{j}]_{j\in J}$ of vectors from $(v_{j})$ are known, the matrix $U_{\alpha}$ is uniquely determined by
$U_{\sigma}[v_{j}\alpha_{j}]_{j\in J}=[v_{\sigma j}]_{j\in J}\hskip 10.00002pt\Longrightarrow\hskip 10.00002ptU_{\sigma}=[v_{\sigma j}]_{j\in J}[v_{j}\alpha_{j}]_{j\in J}^{-1},$
@@
[…]
@@
The action group of the faithful representation of $2\cdot A_{6}$ obtained in Example 5.1 contains $40$ reflections (of order $3$), and it is an irreducible reflection group which appears on the list of [Coh80]. Moreover, the vectors giving the lines are eigenvectors of nontrivial elements of the group, and so the six equiangular lines in $\mathbb{H}^{2}$ can be constructed directly from the reflection group as a group frame (or even from the abstract group $2\cdot A_{6}$).
@@
$|\langle v_{j},v_{k}\rangle|^{2}={1+\langle\vec{v_{j}},\vec{v_{k}}\rangle\over 2}={n-2\over 2(n-1)}\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002pt\langle\vec{v_{j}},\vec{v_{k}}\rangle=-{1\over n-1}.$
@@ 5.1 Concluding remarks
Our focus has been on group frames and equiangular lines. The maximal set of six equiangular lines in $\mathbb{H}^{2}$ comes as the orbit of a quaternionic reflection group, just as the SIC of four equiangular lines in $\mathbb{C}^{2}$ is the orbit of a complex reflection group. However, the known SICs in $\mathbb{C}^{d}$ (with one exception) are orbits of the Weyl-Heisenberg group, which is not a reflection group for $d\geq 3$. The key to constructing quaternionic equiangular lines will be knowing “the right group”. This group might come from numerical constructions, using the techniques of this last section, or from the theory of group representations over $\mathbb{H}$ (which is in its infancy). The construction of sets of tight quaternionic lines may also offer insight into Zauner’s conjecture. Another direction of similar interest is that of optimal packings in quaternionic projective space $\mathbb{H}\mathbb{P}^{k}$.
@@
5.2 Acknowledgements
We wish to thank the referees of an earlier version of this paper who alerted us to the work of [CKM16], which we were unaware of.

### v3

@@ 5 Projective unitary equivalence
This leads to a “linear system” $C(W^{*}W)=(V^{*}V)C$ in the scalars $\alpha_{j}$. However, due to the noncommutativity of the quaternions, this can not be solved by Gauss elimination, unless one first converts it to a linear system over $\mathbb{R}$ (in the coordinates of the $\alpha_{j}$). What is usually done in the real and complex cases is to consider a collection of invariants: the $m$-products, which completely characterise projective unitary equivalence [CW16]. We now look at the analogue of these (also see [KMW25] for subspaces of $\mathbb{C}^{d}$).
@@
$\Delta\bigl{(}(U$ $v_{j_{1}})\alpha_{j_{1}},(Uv_{j_{2}})\alpha_{j_{2}},\ldots,(Uv_{j_{m}})\alpha_{j_{m}}\bigr{)}$
@@
$=\overline{\alpha_{j_{1}}}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})\alpha_{j_{1}}$
$=\alpha_{j_{1}}^{-1}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})\alpha_{j_{1}},$
the $m$-products are projective unitary invariants of $(v_{j})$ up to similarity (congruence), and real frames are characterised by having real $m$-products. Since a quaternion $q$ is determined up to similarity (conjugation in $\mathbb{H}\setminus\{0\}$, which is equivalent to conjugation by unit scalars) by its real part $\mathop{\rm Re}\nolimits(q)$ and its norm $|q|$, we can define (reduced) $m$-products as a pair of real numbers
@@
These are projective unitary invariants. For the complex case, the $m$-products are projective unitary invariants, which depend only on the cycle $(j_{1},\ldots,j_{m})$, and a small set of $m$-products corresponding to a basis for the cycle space of the frame graph of $(v_{j})$ provide a set of invariants which characterise projective unitary equivalence (see [CW16]). We can not yet make a similar claim in the quaternionic case, though we do imagine that the $m$-products do characterise projective unitary equivalence.
The dependence of $m$-products on only the associated $m$-cycle in the frame graph does follow, by the calculation (for nonzero $m$-products)
$a^{-1}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})a=\Delta(v_{j_{2}},v_{j_{3}},\ldots,v_{j_{m}},v_{j_{1}}),\qquad a={\langle v_{j_{1}},v_{j_{2}}\rangle\over|\langle v_{j_{1}},v_{j_{2}}\rangle|},$
@@
$\qquad=|\langle v_{1},v_{2}\rangle|^{2}|\langle v_{2},v_{3}\rangle|^{2}\cdots|\langle v_{k-1},v_{k}\rangle|^{2}\Delta(v_{k},v_{k+1},\ldots,v_{m},v_{1},w_{1},w_{2},\ldots,w_{n}).$
@@
$\langle v_{k},v_{k+1}\rangle\langle v_{k+1},v_{k+2}\rangle\cdots\langle v_{m-1},v_{m}\rangle\langle v_{m},v_{1}\rangle\langle v_{1},v_{2}\rangle\cdots\langle v_{k-2},v_{k-1}\rangle\langle v_{k-1},v_{k}\rangle$
$\qquad\times\langle v_{k},v_{k-1}\rangle\langle v_{k-1},v_{k-2}\rangle\cdots\langle v_{2},v_{1}\rangle\langle v_{1},w_{1}\rangle\langle w_{1},w_{2}\rangle\cdots\langle w_{n-1},w_{n}\rangle\langle w_{n},v_{k}\rangle,$
which simplifies to the right hand side, since $\langle v_{j-1},v_{j}\rangle\langle v_{j},v_{j-1}\rangle=|\langle v_{j-1},v_{j}\rangle|^{2}\in\mathbb{R}$ commutes with any quaternion.
@@
and the permutations of the vectors which preserve these $5$-products is indeed $A_{6}$. Thus the projective symmetry group is a subgroup of $A_{6}$. With the present theory, this does not yet establish that $A_{6}$ is the projective symmetry group.
@@
Once the unit scalars $\alpha_{j}$ corresponding to a basis $v_{j}]_{j\in J}$ of vectors from $(v_{j})$ are known, the matrix $U_{\sigma}$ is uniquely determined by
$U_{\sigma}[v_{j}\alpha_{j}]_{j\in J}=[v_{\sigma j}]_{j\in J}\quad\Longrightarrow\quad U_{\sigma}=[v_{\sigma j}]_{j\in J}[v_{j}\alpha_{j}]_{j\in J}^{-1},$
@@
$a=(12)(34)\quad\hbox{(order $2$)},\qquad b=(1235)(46)\quad\hbox{(order $4$)}$
@@
The action group of the faithful representation of $2\cdot A_{6}$ obtained in Example 5.1 contains $40$ reflections (of order $3$), and it is an irreducible reflection group which appears on the list of [Coh80]. The vectors giving the lines are eigenvectors of nontrivial elements of the group, and so the six equiangular lines in $\mathbb{H}^{2}$ can be constructed directly from the reflection group as a group frame (or even from the abstract group $2\cdot A_{6}$) [Wal24].
@@
$|\langle v_{j},v_{k}\rangle|^{2}={1+\langle\vec{v_{j}},\vec{v_{k}}\rangle\over 2}={n-2\over 2(n-1)}\quad\Longleftrightarrow\quad\langle\vec{v_{j}},\vec{v_{k}}\rangle=-{1\over n-1}.$
@@ 5.1 Concluding remarks
Our focus has been on group frames and equiangular lines. The maximal set of six equiangular lines in $\mathbb{H}^{2}$ comes as the orbit of a quaternionic reflection group, just as the SIC of four equiangular lines in $\mathbb{C}^{2}$ is the orbit of a complex reflection group. However, the known SICs in $\mathbb{C}^{d}$ (with one exception) are orbits of the Weyl-Heisenberg group, which is not a reflection group for $d\geq 3$. The key to constructing quaternionic equiangular lines in this way will be knowing “the right group”. This group might come from numerical constructions, using the techniques of this last section, or from the theory of group representations over $\mathbb{H}$ (which is in its infancy). The construction of sets of tight quaternionic lines may also offer insight into Zauner’s conjecture. Another direction of similar interest is that of optimal packings in quaternionic projective space $\mathbb{H}\mathbb{P}^{k}$.
@@
Conflict of interest statement
The author declared that he has no conflict of interest.